If f(x)=3x-7, and g(x)=-x^2+7x find f[g(3)]
Answer by stanbon(75887)

If f(x)=3x-7, and g(x)=-x^2+7x find f[g(3)]
--------------------------
f[g(3)] = f[-3^2+7*3] = f[-9+21] = f(12) = 3*12-7 = 29
===================
Cheers,
Stan H.
===================